OpenStudy (anonymous):

List two points on the parabola -> x=y^2-2

OpenStudy (anonymous):

here it's best to choose any y-value to get the x value... try y=0....

OpenStudy (anonymous):

Huh? What y values can i choose?

OpenStudy (unklerhaukus):

i would choose ones that fit on your graph, y=0 is a easy choice, and a good choice, but you can choose any real number for y

OpenStudy (anonymous):

i dont understand this at alll. So would (-1,0) be a point?

OpenStudy (unklerhaukus):

\[x=y^2-2\] if you set y=0 \[x=0^2-2\]\[x=-2\] so (x,y) =(-2,0) is a point that satisfies the equation of this parabola

OpenStudy (unklerhaukus):

now set y=1 \[x=1^2-2\]\[x=1-2\]\[x=-1\]so (-1,1) is also a point that satisfies the parabola

OpenStudy (unklerhaukus):

try y=2 what do you get for x?

OpenStudy (anonymous):

x=2^2-2 x=4-2 x=2 (2,2) right?

OpenStudy (unklerhaukus):

right, ! you can try any value for y and you will find a value for x these order pairs (x,y)s are all points on the parabola

OpenStudy (anonymous):

Oh, okay, Thank you so much. That makes since. Can you help me find two point for y=x^2/4

OpenStudy (anonymous):

where would i plug in a number?

OpenStudy (anonymous):

for that one you can choose any x-value...

OpenStudy (anonymous):

So say i chose 4, it would be: 4^2/4 so, 16/4 = 4

OpenStudy (unklerhaukus):

yeah choose x= some number and plug into \[y=\frac{x^2}{4} \rightarrow \text{(some number)}^2/4=\text{some other number}\] and you will find \[(\text{some number} , \text{some other number})\]

OpenStudy (unklerhaukus):

so that corresponds to the point (...,...)

OpenStudy (anonymous):

So (4,4) is a point?

OpenStudy (unklerhaukus):

correct!

OpenStudy (anonymous):

& then (6,9) is a point?

OpenStudy (anonymous):

y=6^2/4 36/4=9 (6,9)

OpenStudy (anonymous):

yes